A stone of mass 59kg released from a height of 2m above the ground. If the stone falls freely to a height of 5m above the ground what is it's velocity
1
Expert's answer
2019-04-18T11:33:44-0400

From the conservation of energy:


"mgH=mgh+0.5mv^2"

"v=\sqrt{2g(H-h)}"

"v=\sqrt{2(9.8)(20-5)}=17\frac{m}{s}"